OpenStudy (anonymous):

can some check cscθ secθ/csc^2θ+sec^2θ csc^2x=1/sin^2x x=1/cosx sec^2x=x1/cosx (1/sin^2x)+(1/cos^2x) how do i simplfy?

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

plz someone

Nnesha (nnesha):

is it one question or 4 ? :o

OpenStudy (anonymous):

1

OpenStudy (anonymous):

i showed my work

OpenStudy (anonymous):

i wanted to know if it is correct

Nnesha (nnesha):

alright let me check it is this ur question \[\large\rm \frac{ \csc \theta \sec \theta }{ \csc^2 \theta } =\sec^2 \theta \]

OpenStudy (anonymous):



Nnesha (nnesha):

ohh okay wait a sec i'll check

OpenStudy (anonymous):

k

Nnesha (nnesha):

i don't get what you mean by this `x=1/cosx ` but looks like you're working on the denominator \[\large\rm \frac{ \csc \theta \sec \theta }{ \csc^2 \theta + \sec^2 \theta } \rightarrow \frac{ \csc \theta \sec \theta }{ \frac{ 1 }{ \sin^2\theta }+\frac{ 1 }{ \cos^2 \theta }}\]

OpenStudy (anonymous):

yes the denominator

Nnesha (nnesha):

alright good so far now first lets deal with the denominator \[\large\rm \frac{ \csc \theta \sec \theta }{ \csc^2 \theta + \sec^2 \theta } \rightarrow \frac{ \csc \theta \sec \theta }{ \color{Red}{\frac{ 1 }{ \sin^2\theta }+\frac{ 1 }{ \cos^2 \theta } }}\] find common denominator

OpenStudy (anonymous):

cos

Nnesha (nnesha):

\[\large\rm \frac{ 1 }{ \sin^2\theta }+\frac{ 1 }{ \cos^2 \theta }\] hmm no both denominators are different so we should multiply them

OpenStudy (anonymous):

csc2+sec2

Nnesha (nnesha):

here is an example when we find common denominators \[\large\rm \frac{ 1 }{ x }+\frac{1}{y}\] both denominators aren't the same so we should multiply them \[\frac{ y +x}{ xy }\] and when we find common denominator we should multiply numerator of first fraction by the denominator of 2nd fraction multiply numerator of 2nd fraction by the denominator of first fraction

Nnesha (nnesha):

no we changed csc^2+sec^2 to 1/sin^2theta +1/cos^2

Nnesha (nnesha):

\[\large\rm \frac{ 1 }{ \sin^2\theta }+\frac{ 1 }{ \cos^2 \theta }\] what's common denomiantor ?

OpenStudy (anonymous):

2

Nnesha (nnesha):

hmm there isn't any 2 let's say sin^2 theta = a and cos^2 theta = b \[\large\rm \frac{ 1 }{a }+\frac{ 1 }{b}\] what's the common denominator ?

OpenStudy (anonymous):

cos

Nnesha (nnesha):

not just cos hmm look at the example i gave you

OpenStudy (anonymous):

srry sin would be considered

Nnesha (nnesha):

yes both denominators aren't the same so we should multiply them and btw it's sin^2 and cos^2 not just sin cos

Nnesha (nnesha):

\[\frac{ ??? +??? }{ \sin^2 \theta \cos^2\theta }\] what would you get at the numerator ?

OpenStudy (anonymous):

it would be csctheta sectheta

Nnesha (nnesha):

hmm no that's the numerator but i was talking about the numerator of bottom fraction \[\large\rm \frac{ \csc \theta \sec \theta }{ \color{Red}{\frac{ 1 }{ \sin^2\theta }+\frac{ 1 }{ \cos^2 \theta } }}\rightarrow \frac{ csc \theta sec \theta }{\frac{???+??}{sin^2 \theta +cos^2 \theta }}\]

OpenStudy (anonymous):

1/sinx 1/cos

Nnesha (nnesha):

here is an example when we find common denominators \[\large\rm \frac{ 1 }{ x }+\frac{1}{y}\] both denominators aren't the same so we should multiply them \[\frac{ y +x}{ xy }\] and when we find common denominator we should multiply numerator of first fraction by the denominator of 2nd fraction multiply numerator of 2nd fraction by the denominator of first fraction

Nnesha (nnesha):

` multiply numerator of first fraction by the denominator of 2nd fraction` `multiply numerator of 2nd fraction by the denominator of first fraction `

OpenStudy (anonymous):

so it would be cos theta and sin theta

Nnesha (nnesha):

hmm not just cos it's cos^2

OpenStudy (anonymous):

ok

Nnesha (nnesha):

don't forget the sign hmm no that's the numerator \[\huge\rm \ \rm \frac{ csc \theta sec \theta }{\frac{cos^2 \theta +sin^2 theta }{sin^2 \theta *cos^2 \theta }}\] i made a typo earlier it's sin^2 times cos^2 not +

Nnesha (nnesha):

now we can rewrite csc and sec in terms of cos sin \[\huge\rm \frac{ \frac{ 1 }{ \sin \theta } *\frac{1}{\cos^2\theta }}{ \frac{ \cos^2 \theta +\sin^2 \theta }{\sin^2 \theta *\cos^2 \theta}}\]

Nnesha (nnesha):

tr to simplify that remember the special identity \[\sin^2 \theta +\cos^2 \theta =1\]tr to simplify this remember the special identity \[\sin^2 \theta +\cos^2 \theta =1\]

Nnesha (nnesha):

try*

OpenStudy (anonymous):

one sec

Nnesha (nnesha):

sure take u time :=)

Nnesha (nnesha):

ur**

OpenStudy (anonymous):

sintheta

Nnesha (nnesha):

wait there is a typo

OpenStudy (anonymous):

nnesha could you please help me on the question I tagged you in if you can?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha now we can rewrite csc and sec in terms of cos sin \[\huge\rm \frac{ \frac{ 1 }{ \sin \theta } *\frac{1}{\cos^2\theta }}{ \frac{ \cos^2 \theta +\sin^2 \theta }{\sin^2 \theta *\cos^2 \theta}}\] \(\color{blue}{\text{End of Quote}}\) it's\[\huge\rm \frac{ \frac{ 1 }{ \sin \theta } *\frac{1}{\color{Red}{\cos\theta} }}{ \frac{ \cos^2 \theta +\sin^2 \theta }{\sin^2 \theta *\cos^2 \theta}}\] cos theta

Nnesha (nnesha):

no it 's not sin please show the steps how did you get that

OpenStudy (anonymous):

so cos theta is the answer

Nnesha (nnesha):

how did you get tthat ?

OpenStudy (anonymous):

ok so sin theta was right

Nnesha (nnesha):

how did you get that please show ur work

OpenStudy (anonymous):

i guessed it

Nnesha (nnesha):

after doing all this work now we should guess the answer ???

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

i also have notes

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ 1 }{ \sin \theta } *\frac{1}{\cos\theta }}{ \frac{\color{Red}{ \cos^2 \theta +\sin^2 \theta} }{\sin^2 \theta *\cos^2 \theta}}\] cos^2 x+sin^2 x= ??

OpenStudy (anonymous):

1

Nnesha (nnesha):

right \[\huge\rm \frac{ \frac{ 1 }{ \sin \theta } *\frac{1}{\cos\theta }}{ \frac{\color{Red}{ 1} }{\sin^2 \theta *\cos^2 \theta}}\] now change division to multiplication

Nnesha (nnesha):

\[\frac{ 1 }{ \sin \theta }*\frac{1}{\cos \theta} = \frac{1}{ \sin \theta*\cos \theta}\]

OpenStudy (anonymous):

thats my answer

Nnesha (nnesha):

no.

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ 1 }{ \sin \theta*cos \theta } }{ \frac{\color{Red}{ 1} }{\sin^2 \theta *\cos^2 \theta}}\] now change division to multiplication

OpenStudy (anonymous):

csc3(θ)sec3(θ)

Nnesha (nnesha):

hmm here is an example \[\large\rm \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=\frac{ a }{ b }*\frac{ d }{ c }\] multiply top fraction by the `reciprocal ` of the bottom fraction

OpenStudy (anonymous):

sine

Nnesha (nnesha):

show me how did you get that ?

OpenStudy (anonymous):

i did 1/sintheta*costheta

Nnesha (nnesha):

tag my name ilost ur post

Nnesha (nnesha):

nvm found it

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

alright what's the reciprocal of bottom fraction ??

OpenStudy (anonymous):

Would it be cos sin

OpenStudy (anonymous):

I suck at math

Nnesha (nnesha):

\[\huge\rm \frac{ \color{Red}{\frac{ 1 }{ \sin \theta*cos \theta } }}{ \color{blue}{\frac{ 1 }{\sin^2 \theta *\cos^2 \theta}}}\] what's the reciprocal of bottom fraction ?

Nnesha (nnesha):

this is like pre-calculus but we are dealing with algebra stuff practice will make it easy

Nnesha (nnesha):

here is an example reciprocal of x/y is y/x

OpenStudy (anonymous):

Cos theta/sin theta

Nnesha (nnesha):

hmm no \[\frac{ 1 }{ \sin^2 \theta* \cos^2 \theta }\] just flip the fraction !

OpenStudy (anonymous):

Sin2theta cos2theta/1

Nnesha (nnesha):

\[\frac{ 1 }{ \cos \theta*\sin \theta } *\frac{\sin^2 \theta \cos^2 \theta}{1}\] now simplify please

OpenStudy (anonymous):

Cos theta sin theta

Nnesha (nnesha):

correct

OpenStudy (anonymous):

So that's my answer

Nnesha (nnesha):

yep.

OpenStudy (anonymous):

Thank you for the help I really do appreciate it:)

Nnesha (nnesha):

np